OpenStudy (lifeisadangerousgame):

I need to know how to do this please Which equation is in standard form? Answer 4x = y + 12 y = 3x - 18 8x + 10y = 21 3x + 9y + 9 = 4

OpenStudy (anonymous):

the third one:8x + 10y = 21 because its of the form ax+by=c

OpenStudy (lifeisadangerousgame):

ax + by = c is standard form?

OpenStudy (anonymous):

and y = 3x - 18 is in slope intercept form y=mx+c

OpenStudy (anonymous):

yes standard form of a st. line is ax+by+c=0, so the eq. 8x + 10y = 21 can be taken as 8x+10y-21=0

OpenStudy (lifeisadangerousgame):

Ok! Thank you :)

OpenStudy (anonymous):

i thought u had more questions, when you typed

OpenStudy (anonymous):

1. Ax + By = C 2. A, B, C are integers (positive or negative whole numbers) 3. No fractions nor decimals in standard form. 4. Traditionally the "Ax" term is positive. http://www.algebralab.org/studyaids/studyaid.aspx?file=algebra1_5-5.xml

OpenStudy (anonymous):

gud add up @marinaDL

OpenStudy (anonymous):

I like be detailed.